# A Global Minimum Tax for Large Firms Only:

### v5

@@ A.3. Proofs of Propositions 3 and 4
$\Delta\ \begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}\\ <0&\text{if}\ t_{M}>\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}\end{cases},$
@@
The marginal effect of the GMT is
$\frac{\partial G_{n}^{1}}{\partial t_{M}}=\frac{H\lambda^{2}t_{M}+\delta(\lambda-1)^{2}}{\delta(2\lambda-1)}\Pi>0,\qquad\frac{\partial G_{h}^{1}}{\partial t_{M}}=\frac{\lambda[\delta(\lambda-1)-2H\lambda t_{M}]}{\delta(2\lambda-1)}\Pi\begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)}{2H\lambda}\\ <0&\text{if}\ t_{M}>\frac{\delta(\lambda-1)}{2H\lambda}\end{cases},$
@@
In Regime 1, the welfare levels of the non-haven and the world rise after the GMT introduction and they increase along with a gradual increase in the GMT rate. On the other hand, the welfare level of the haven rises if $t_{M}<\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$ after the GMT introduction and it falls if $t_{M}>\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$.
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
For the non-haven country, we see
@@
In Regime 2, the welfare levels of the non-haven and the world therefore rise after the GMT introduction and they increase along with a gradual increase in the GMT rate. On the other hand, the welfare level of the haven declines after the GMT introduction and it is independent of $t_{M}$.
Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}=\frac{2\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
For the non-haven country, we have
@@
In Regime 3, the welfare levels of the non-haven, the haven, and the world all rise after the GMT introduction and they increase along with a gradual increase in the GMT rate.
@@
In Regime 4, the welfare levels of the non-haven and the world rise after the GMT introduction and they increase along with a gradual increase in the GMT rate. On the other hand, the welfare level of the haven declines after the GMT introduction and it is independent of $t_{M}$.
This proves Proposition 3 and a part of Proposition 4. The proof of the results on the regime-switching points is given in Appendix 3.
A.4. Regime-switching points
@@
This completes the proof of Proposition 4.
A.5. Proof of Proposition 5
The effects of an increase in the GMT coverage rate, $\phi$, on welfare in the non-haven, the haven, and the world are given as follows.
@@ A.5. Proof of Proposition 5
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
For the non-haven country, we see
@@
The sign of the derivative is determined by that of $\Theta_{n}$. From $\Theta_{n}(t_{M}=t_{M}^{1})>0$, $\Theta_{n}(t_{M}=t_{M}^{1})>0$, and the fact that $\Theta_{n}(t_{M})$ is linear in $t_{M}$, it follows that $\Theta_{n}>0$.
For the haven country, we see
@@
The sign of the derivative is determined by that of $\Theta_{h}$. From $\Theta_{h}(t_{M}=t_{M}^{1})<0$, $\Theta_{h}(t_{M}=t_{M}^{1})<0$, and the linearity of $\Theta_{h}(t_{M})$ in $t_{M}$, it follows that $\Theta_{h}<0$.
@@
The sign of the derivative is determined by that of $\Theta_{W}$. From $\Theta_{W}(t_{M}=t_{M}^{1})>0$, $\Theta_{W}(t_{M}=t_{M}^{1})>0$, and the linearity of $\Theta_{W}(t_{M})$ in $t_{M}$, it follows that $\Theta_{W}>0$.
Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$. We see
$\frac{\partial G_{n}^{3}}{\partial\phi}=\frac{H(4\lambda-1)t_{M}^{2}}{8\delta}\Pi>0,\quad\frac{\partial G_{h}^{3}}{\partial\phi}=-\frac{H\lambda t_{M}^{2}}{8\delta}\Pi<0,\quad\frac{\partial G_{W}^{3}}{\partial\phi}=\frac{H(2\lambda-1)t_{M}^{2}}{4\delta}\Pi>0,$
@@
This completes the proof of Proposition 5.
Appendix B: Data for Figure 1
We select MNEs from the Orbis based on the following criteria.
The numbers on which Figure 1 is based are given in Table B1. In column A, to calculate the pre-tax profits of all MNEs in a given year, we exclude those with missing revenues (“Operating revenue (Turnover)”) and pre-tax profits (“P/L before tax”). In column B, MNEs subject to the GMT are those with annual revenues of no less than 750 million EUR in at least two years of the last four years. Using column A and B, column C reports the share of MNEs subject to the GMT in terms of pre-tax profits and numbers.
Table B1: MNEs in the Orbis database
A: All MNEs B: MNEs $\geq$ 750mEUR C: Share $=$ B/A
2018 2320 8656 2066 2333 0.89 0.27
2019 2430 8416 2216 2437 0.91 0.29
2020 1985 7920 1793 2228 0.90 0.28
2021 3564 8803 3239 2627 0.91 0.30
Source: Orbis database, own calculations.
Note: Pre-tax profits are in billion EUR.

### v6

@@ Appendix C Welfare Effects of the Introduction of GMT
$\Delta\ \begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}\equiv t_{M}^{+}\\ <0&\text{if}\ t_{M}>t_{M}^{+}\end{cases},$
@@
In Regime 1, the welfare effect of the introduction a GMT rate is positive in the non-haven and the world. In the haven, it is positive if $t_{M}\in[t_{M}^{0},t_{M}^{+})$ and negative if $t_{M}\in(t_{M}^{+},t_{M}^{1}]$.
A marginal increase in $t_{M}$ leads to
$\frac{\partial G_{n}^{1}}{\partial t_{M}}=\frac{H\lambda^{2}t_{M}+\delta(\lambda-1)^{2}}{\delta(2\lambda-1)}\Pi>0,$
$\frac{\partial G_{h}^{1}}{\partial t_{M}}=\frac{\lambda[\delta(\lambda-1)-2H\lambda t_{M}]}{\delta(2\lambda-1)}\Pi\begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)}{2H\lambda}\equiv t_{M}^{++}\\ <0&\text{if}\ t_{M}>t_{M}^{++}\end{cases},$
@@
In Regime 1, the welfare effect of a marginal increase in the GMT rate is positive in the non-haven and the world. In the haven, it is positive if $t_{M}\in[t_{M}^{0},\delta(\lambda-1)/(2H\lambda))$ and negative if $t_{M}\in(\delta(\lambda-1)/(2H\lambda),t_{M}^{1}]$.
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$.
For the non-haven, we see
@@
In Regime 2, both the introduction and a marginal increase of the GMT rate raise welfare in the non-haven and the world, but reduce it in the haven.
Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}=\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
For the non-haven, we have
@@
In Regime 3, the introduction of a GMT rate increases welfare in the non-haven and the world, while reducing it in the haven. A marginal increase in the GMT rate raises welfare in the non-haven, the haven and the world.
@@
In Regime 4, the introduction of a GMT rate increases welfare in the non-haven and the world, while reducing it in the haven. A marginal increase in the GMT rate raises welfare in the non-haven and the world, but has no effect on welfare in the haven. ∎
Appendix D Welfare Effects of Increasing the GMT Rate
The following proposition characterises the welfare effects of a marginal increase in the GMT rate $t_{M}$.
Proposition D.1
(Increase in the GMT rate $t_{M}$)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text. A marginal increase in the GMT rate has the following effects:
  • (i)
    welfare in the non-haven increases in all regimes, except at the regime-switching point $t_{M}^{1}$, where it falls discontinuously.
  • (ii)
    welfare in the (representative) haven increases in Regime 1 as long as $t_{M}<t_{M}^{++}\equiv\frac{\delta(\lambda-1)}{2H\lambda}$, then decreases thereafter in Regime 1 and in Regime 2. It increases in Regime 3, falls discontinuously at the regime-switching point $t_{M}^{3}$, and remains unchanged in Regime 4.
  • (iii)
    welfare in the world increases in all regimes, except at the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, where it falls discontinuously.
As is true for the introduction of the GMT (Proposition C.1 in Section C of this Supplemental Appendix), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium. An exception is the regime-switching point $t_{M}^{1}$ where the havens being to split their tax rates and set a discretely lower tax rate for small MNEs. Small MNEs then shift more profits to the havens, hurting the non-haven.
For havens, the welfare (i.e., tax revenue) effects of an increase in the GMT rate reflect two opposing forces: a negative effect from reduced profit-shifting opportunities and a positive effect from a higher tax rate in the non-haven. These opposing forces generate a non-monotonic welfare response. In Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{1}]$), the positive effect dominates initially (when $t_{M}<t_{M}^{++}$), but as the GMT rate rises further, the negative effect prevails (when $t_{M}\geq t_{M}^{++}$). In Regime 2 ($t_{M}\in(t_{M}^{1},t_{M}^{2}]$), the negative effect dominates throughout. In Regime 3 ($t_{M}\in(t_{M}^{2},t_{M}^{3}]$), the positive effect of a higher tax rate dominates again, as the non-haven commits to a uniform GMT rate in order to raise the havens’ tax rates. This regime, where commitment induces effective tax coordination, is similar to the first part of Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{++})$). It ends at the regime-switching point $t_{M}^{3}$ at which the non-haven starts splitting its tax rate. This leads to a discontinuous fall in the non-haven’s tax rate on small MNEs, which must hurt the tax havens. In Regime 4 ($t_{M}\in(t_{M}^{3},1]$), both the non-haven and the havens apply the same GMT rate to large MNEs, preventing them from shifting profits. Consequently, changes in the GMT rate no longer affect the havens’ welfare.
The marginal effects of $t_{M}$ on world welfare are dominated by the effects on the non-haven, and are therefore generally positive. At the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, however, one of the countries begins to split its tax rate and discontinuously lowers the tax rate on small MNEs. Therefore, tax competition is tightened and world welfare falls at the margin.
In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
Proof:  In the proof of Proposition C.1 in Section C, we already showed the welfare effects of a marginal increase in the GMT rate within each regime. Here we show the welfare effects at the regime switching points.
@@ Appendix D Welfare Effects of Increasing the GMT Rate
This completes the proof of Proposition D.1. ∎
Appendix E Welfare Effects of Increasing the GMT Coverage Rate
The following proposition characterizes the welfare effects of a marginal increase in the GMT coverage rate $\phi$.
Proposition E.1
(Increasing in the GMT coverage rate $\phi$)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text. Except for special cases where a regime switch occurs, a marginal increase in the GMT coverage rate of $\phi\in(0,1)$ has the following effects:
  • (i)
  • (ii)
    welfare in the (representative) haven is unaffected in Regime 1 and decreases in Regimes 2 to 4.
  • (iii)
    world welfare is unaffected in Regime 1 and increases in Regimes 2 to 4.
In Regime 1, an increase in $\phi$ has no effect on welfare in any country, because all countries levy uniform tax rates. Changing $\phi$ does have effects, however, in the other regimes, where one or both countries split their tax rates. Since the equilibrium tax differential is always greater for small MNEs than for large MNEs, a higher $\phi$, which implies more large MNEs subject to the GMT, reduces aggregate profit shifting. This unambiguously hurts the haven and benefits the non-haven. From the global perspective, a higher GMT coverage is always desirable, as it reduces pressure on tax competition for the profits of small MNEs.
There are two exceptional cases where increasing $\phi$ causes a regime switch. The first is at $t_{M}=t_{M}^{1}+\epsilon$ with $\epsilon$ being a small positive number, where a rise in $\phi$ pushes Regime 2 back to Regime 1 by weakening the havens’ incentive to split their tax rates. This regime switch benefits the non-haven and the world, while it has no effect on the havens. The other is at $t_{M}=t_{M}^{2}+\epsilon$, where a higher $\phi$ triggers a switch from Regime 3 to Regime 2, inducing the non-haven to set a tax rate above the GMT. This latter regime switch, however, leaves welfare in all countries and the world unaffected.
In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
Proof:  The effects of an increase in the GMT coverage rate, $\phi$, on welfare in the non-haven, the haven, and the world are given as follows.
@@ Appendix E Welfare Effects of Increasing the GMT Coverage Rate
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$.
For the non-haven, we see
@@
The sign of the derivative is determined by that of $\Theta_{n}$. From $\Theta_{n}(t_{M}=t_{M}^{1})>0$, $\Theta_{n}(t_{M}=t_{M}^{2})>0$, and the fact that $\Theta_{n}(t_{M})$ is linear in $t_{M}$, it follows that $\Theta_{n}>0$.
For the haven, we see
@@
The sign of the derivative is determined by that of $\Theta_{h}$. From $\Theta_{h}(t_{M}=t_{M}^{1})<0$, $\Theta_{h}(t_{M}=t_{M}^{2})<0$, and the linearity of $\Theta_{h}(t_{M})$ in $t_{M}$, it follows that $\Theta_{h}<0$.
@@
The sign of the derivative is determined by that of $\Theta_{W}$. From $\Theta_{W}(t_{M}=t_{M}^{1})>0$, $\Theta_{W}(t_{M}=t_{M}^{2})>0$, and the linearity of $\Theta_{W}(t_{M})$ in $t_{M}$, it follows that $\Theta_{W}>0$.
We see
$\frac{\partial G_{n}^{3}}{\partial\phi}=\frac{H(4\lambda-1)t_{M}^{2}}{8\delta}\Pi>0,\qquad\frac{\partial G_{h}^{3}}{\partial\phi}=-\frac{H\lambda t_{M}^{2}}{8\delta}\Pi<0,\qquad\frac{\partial G_{W}^{3}}{\partial\phi}=\frac{H(2\lambda-1)t_{M}^{2}}{4\delta}\Pi>0,$
@@
This completes the proof of Proposition E.1. ∎
Appendix F Supplementary Figures
Figure F.1 shows the welfare levels of the non-haven (panel (a)) and the sum of the havens (panel (b)), and the shifted profits for different GMT rates $t_{M}$.
The tax-competition equilibrium outcomes under partial GMT coverage (solid curves) are compared with those under full GMT coverage (dashed curves). The equilibrium tax rates under full GMT are as follows.
Regime 0: $t_{M}\in\left[0,t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)}\right)$.  Each of the non-haven and the haven sets its own single non-GMT rate:
$t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
Regime 1f: $t_{M}\in\left[t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)},t_{M}^{f}\equiv\frac{\delta(\lambda-1)}{H\lambda}\right]$.  The non-haven chooses a single non-GMT rate, and the (representative) haven chooses a single GMT rate:
$t_{n}^{1f}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{1f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
Regime 2f: $t_{M}\in(t_{M}^{f},1]$.  Both the non-haven and the haven choose a single GMT rate:
$t_{n}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
Let $G_{n}^{f}$ and $G_{h}^{f}$ be the equilibrium payoffs of the non-haven and the haven under full GMT. In Regime 0 ($t_{M}\in[0,t_{M}^{0})$), the equilibrium tax rates are identical under full and partial GMT coverage, as neither country is constrained by the GMT.
We note $t_{M}^{f}>t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$, given $\phi\in(0,1)$ and $\lambda>1$. That is, the equilibrium outcomes under partial GMT begins to diverge from those under full GMT once $t_{M}$ enters the range of Regime 1. This is particularly evident in panels (a) and (c).
Fig. F.1. Equilibrium Welfare and Shifted Profits.
Notes: The figure shows the equilibrium levels of welfare in the non-haven ($G_{n}$) in panel (a) and the sum of the $H=40$ havens ($\sum_{h=1}^{H}G_{h}$) in (b), and the aggregate shifted profits ($\sum_{h=1}^{H}\theta_{h}\Pi$) in (c), for different GMT rates $t_{M}$. The solid curves in each panel are the equilibrium values under the partial GMT ($\phi\in(3/4,1)$); the dashed curves are those under the full GMT ($\phi=1$); horizontal lines are those under no GMT (Regime 0). While the levels of welfare, shifted profits and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.